A ring of mean diameter 30 cm is wound with 200 turns of copper wire carrying a current of 2 

A. The x-section of the magnetic material of the ring is 12 cm2 and its relative permeability is 

1,000. Determine theflux through it.


1
Expert's answer
2021-06-15T06:58:01-0400

Diameter(D)=30cm \\


Circumference=\pi\times D \\


C=30\pi cm \\


Turns (N)=200 \\


Current (I)= 2A \\


Relative permeability (\mu_r)=1000 \\


Cross section area (A)=12cm^2 \\


\mu_o= 4\pi \times 10^{-7} N/A^2 \\


\Phi= \frac {\mu_o\times \mu_r A\times N \times I}{C} \\


\Phi= \frac{4\pi \times 10^{-7} \times 1000 \times 12 \times 200 \times 2}{30\pi} \\


\Phi= 32mWb